OpenStudy (anonymous):

Use the definitions sinθ=y/r, cosθ=x/r and/or tanθ=y/x to prove that cotθsecθ=cscθ

OpenStudy (anonymous):

cot theta = 1/tan theta = x/y Sec theta = 1/Cos theta = r/x Cosec Theta = 1/Sin Theta = r/y Put them in and you'll get the proof..